OpenStudy (daenio):

2. Tristan and Laura were investigating ancient Voodoo beliefs of Mount Hope and Ancaster. In ancient times, the villagers believed that the Voodoo Gods must be happy to have a successful harvest each year. If the probability that the Voodoo Gods would be happy was , a) Show the binomial probability distribution that there would be a successful harvest four years in a row. [5] b) What is the probability the Voodoo Gods will be happy for at least two of the four years? [2]

OpenStudy (daenio):

@jim_thompson5910 @satellite73 @LagrangeSon678 @Mertsj @kropot72 @UnkleRhaukus @radar @Chlorophyll

jimthompson5910 (jim_thompson5910):

You're missing the probability

OpenStudy (daenio):

Tristan and Laura were investigating ancient Voodoo beliefs of Mount Hope and Ancaster. In ancient times, the villagers believed that the Voodoo Gods must be happy to have a successful harvest each year. If the probability that the Voodoo Gods would be happy was 1/5, a) Show the binomial probability distribution that there would be a successful harvest four years in a row. [5] Sorry about that! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

a) If you're only taking 4 years into account and that's it, then the probability is simply (1/5)^4 = 1/625 = 0.0016 ------------------ b) P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 2) = (4 C 2)*(0.2)^(2)*(1-0.2)^(4-2) P(X = 2) = (4 C 2)*(0.2)^(2)*(0.8)^(4-2) P(X = 2) = (6)*(0.2)^(2)*(0.8)^2 P(X = 2) = (6)*(0.04)*(0.64) P(X = 2) = 0.1536 P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 3) = (4 C 3)*(0.2)^(3)*(1-0.2)^(4-3) P(X = 3) = (4 C 3)*(0.2)^(3)*(0.8)^(4-3) P(X = 3) = (4)*(0.2)^(3)*(0.8)^1 P(X = 3) = (4)*(0.008)*(0.8) P(X = 3) = 0.0256 P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 4) = (4 C 4)*(0.2)^(4)*(1-0.2)^(4-4) P(X = 4) = (4 C 4)*(0.2)^(4)*(0.8)^(4-4) P(X = 4) = (1)*(0.2)^(4)*(0.8)^0 P(X = 4) = (1)*(0.0016)*(1) P(X = 4) = 0.0016 So P(X >= 2) = 0.1536+0.0256+0.0016 P(X >= 2) = 0.1808 which means that the probability the Voodoo Gods will be happy for at least two of the four years is 0.1808

OpenStudy (daenio):

Wow man. I didn't think it involved that much work. Thank you very much for taking the time for a detailed answer! :)

jimthompson5910 (jim_thompson5910):

you're welcome